If 2=10^a and 3=10^b, solve for x in the following, giving your answers in terms of a and b.
a) 10^x=6
b) 10^x=24

Answer by Alan3354(69443)

If 2=10^a and 3=10^b, solve for x in the following, giving your answers in terms of a and b.
a) 10^x=6
6 = 2*3 = 10^a*10^b = 10^(a+b)
x = a+b
=====================
b) 10^x=24
Similar to the 1st one